OpenStudy (anonymous):

HELP ME SOLVE FOR x: −2x + 5 < 7

OpenStudy (anonymous):

You treat the < as an equal sign.

OpenStudy (scarlettred):

-2x + 5 = 7 -2x = 7 - 5 -2x = 2 x = -1

OpenStudy (anonymous):

< or >

OpenStudy (anonymous):

the sign didn't change. So it would be <